OpenStudy (cicilybailey14):

Brainteaser :) 2+3=8, 3+7=27, 4+5=32, 5+8=60, 6+7=72, 7+8=?? Solve it?

TheSmartOne (thesmartone):

2+3=5 a+b=c 2+3=8 a+b=c*a-a 2+3=5*2-2 2+3=10-2 2+3=8 So 7+8=15 7+8=15*7-7 7+8=105-7 7+8=98

TheSmartOne (thesmartone):

@Cicilybailey14

OpenStudy (cicilybailey14):

oh thanks...haha

TheSmartOne (thesmartone):

:) :D